# The Challenges of Detecting Gases in Exoplanet Atmospheres

### v1

@@ Abstract
Claims of detections of gases in exoplanet atmospheres often rely on comparisons between models including and excluding specific chemical species. However, the space of molecular combinations available for model construction is vast and highly degenerate. Only a limited subset of these combinations is typically explored for any given detection. As a result, apparent detections of trace gases risk being artifacts of incomplete modeling rather than robust identification of atmospheric constituents, especially in the low signal-to-noise regime. We illustrate these challenges using the sub-Neptune K2-18 b, where recent claims of a potential biosignature detection vanish when the considered model space is expanded. We show that numerous alternative models without potential biosignature gases provide equivalent or better fits to the observations. We demonstrate that the significance of a claimed detection relies on the choice of models being compared, and that model preference does not necessarily imply the presence of any specific gas.
@@
While the atmospheric retrieval approach is generally regarded as agnostic and data-driven, it still has the potential to mislead regarding the strength of an apparent detection, and the subsequent interpretation of planetary conditions. The field has adopted Bayesian model comparison as a standard method to determine the presence of chemical species and physical processes in exoplanet atmospheres[2, 3, 4]. However, concerns have emerged that relying on the Bayesian evidence as the sole goodness-of-fit metric is insufficient for robust atmospheric interpretation[5]. Recent work has shown the advantages of using methods beyond the Bayesian evidence that provide insight into which parts of a dataset drive inferences[6, 7], and of averaging inferences over models instead of selecting between them[8]. While these studies highlight important limitations, the broader impact of these issues — especially in regimes lacking strong spectral features that can be clearly identified, attributed, and interpreted (see Key Criteria from ref.[9]) — remains to be explored.
Here we demonstrate that Bayesian model comparison does not always translate into robust detections of atmospheric constituents. Pairwise model comparisons only quantify the relative fit to the data of one model against another. When all considered candidate models are poor representations of reality, the best-performing model is simply the least inadequate and may not necessarily lead to reliable interpretations of the data. Conversely, when all candidate models adequately fit a spectrum, a preference for one model over another does not rule out the worse-performing model. We highlight scenarios in which reliance on Bayesian evidence alone, coupled with exploration of only a narrow part of the model space, has led to contradictory interpretations. We demonstrate how these issues manifest in a case study of the sub-Neptune K2-18 b[10, 11], where the inferred atmospheric properties can be shaped as much by preconceived notions for what this planet ought to be like, as by the observational data.
@@
Determining appropriate, data-driven model assumptions was an early challenge in analyzing exoplanet spectra given our limited a priori knowledge of these atmospheres. Bayesian model comparison, and in particular the use of Bayes factors, was introduced to compare competing hypotheses in exoplanetary atmospheric studies[2]. Ref.[2] introduced an explicit methodology of computing the Bayesian evidence for one retrieval model that should cover the full prior hypothesis space, and comparing it to retrieval models for which one individual molecular species (or type of aerosol) was removed at a time from that otherwise full prior hypothesis space. Although exploring a complete prior hypothesis space may not be possible, as explained below, this remove-one-at-a-time approach was intended to ensure that a high confidence in the presence of a particular atmospheric constituent was reported only if no other reasonable constituent could have plausibly resulted in the observed data (Section 2.2.2 in ref.[2]). In the years that followed, this approach of comparing Bayesian evidence was quickly adopted by the field, and became popular for its promise of applying “Occam’s razor” to penalize the fit of more complex models over simpler ones that were not sufficiently supported by the data. The availability of efficient computational tools for parameter estimation and evidence calculation for a pair of given models[12] was also critical for its wide adoption.
@@
Given the practical impossibility of constructing a full prior hypothesis space, and no well-established conventions in the field, there is a risk of Bayesian model comparison being used to evaluate the relative preference between retrieval models even when neither adequately represents the full prior hypothesis space. Detections could then be reported based on Bayes factors between models that each include only a limited subset of chemically plausible species, without comparison or contextualization relative to a more comprehensive hypothesis space. A particularly dangerous scenario would be that in which two similarly incomplete or inadequate models are compared, since there is a risk of interpreting any model preference as a detection. Even more troublesome would be to directly convert this model preference into a “detection significance,” expressed in terms of [13], losing the relative nature of this metric and further contributing to misinterpretation.
@@
Use of model comparisons in isolation, without acknowledgment of the limitations in our models, can contribute to an increasingly opaque definition of detection. Any detection significance is computed relative to a chosen reference model and therefore strongly depends on the included species and parameters. Statements of model preference without qualification of the reference model are arbitrary. Furthermore, because the number of plausible molecular species is large — hundreds are available via the HITRAN database[14, 15] for example — meaningfully exploring any significant fraction of the hypothesis model space is computationally unfeasible. This practical limitation is an invitation to recognize the subjective nature in building an appropriate model space. Take for instance Figure 1, which shows how the number of unique models and pairwise model comparisons needed to exhaustively explore the detection space grows with the number of molecules considered. Even for a modest set of molecular species, the number of unique composition models for an atmosphere is , and the number of pairwise model comparisons needed to fully explore the detection space is . While a complete exploration of model space is not currently practical, any detection claimed through pairwise model comparison should at minimum be contextualized within the broader set of plausible comparisons that could have yielded equal or greater significance. Credible detection claims demand systematic elimination of credible alternatives.
@@
JWST has delivered spectra of some of the smallest and coldest exoplanets studied to date[16, 17, 18, 19, 20, 21, 22, 23, 24]. Several of these observations remain statistically consistent with a featureless spectrum despite the improved precision (see Fig. 2). Bayesian model comparison quantifies the relative preference between two models. Frequentist approaches alternatively test whether a null hypothesis, such as the absence of an atmosphere, can be rejected at a given confidence level (see Methods). This distinction is crucial when interpreting atmospheric retrieval results, particularly as Bayes factors are often mapped to frequentist significances using calibration tables[2, 13], risking a blurring of their fundamentally different meanings. Bayesian evidence statements should always be understood as relative preferences between models and not absolute measures of detection probability.
Degeneracies between atmospheric parameters contribute to the complexity of exoplanet atmospheric retrieval. Numerous molecules have overlapping absorption bands, meaning that several chemical species can often provide comparably good explanations for a given portion of an observed spectrum. A clear example of this degeneracy appears in the analysis of the –m HST/WFC3 transmission spectrum of the sub-Neptune K2-18 b[10]. The spectrum exhibited a prominent feature near m, initially attributed to H2O at [10], [25], and [26, 27] significance by four independent analyses, all using Bayesian model comparisons. If interpreted as a frequentist significance, this would imply a 99.7% probability that H2O is present. However, a subsequent study demonstrated that model realizations with methane as the dominant absorber instead of water also provided adequate fits to the data ()[28].
Follow-up observations with JWST/NIRISS and NIRSpec, spanning a broader spectral range (0.9–5.2m), revealed a preference for methane absorption at – significance, depending on model assumptions, with no strong evidence for H2O features[22, 29]. The extended coverage from JWST helped resolve the degeneracy between H2O and CH4 at HST/WFC3 wavelengths. Comparing the HST and JWST observations, it appears that some previous attributions to H2O may have been influenced by two data points near the blue edge of the WFC3 bandpass (see Figure 6 in ref.[22]). Furthermore, one earlier study did not consider CH4 absorption in its atmospheric models[25]. Overall, the case of the HST/WFC3 spectrum of K2-18 b highlights the risks of interpreting model preference as a definitive chemical detection—particularly when the data are limited in spectral coverage, the models are highly simplified, the data is at a threshold signal-to-noise level for detecting features, and degeneracies remain unresolved.
We note that given the limited availability of telescope resources, JWST observations are frequently optimized to yield 3 detections, especially for small exoplanets. This strategy maximizes the number of targets observed but necessarily results in spectra that operate near the threshold of signal detectability. By contrast, many of these challenges are mitigated in higher signal-to-noise regimes. For example, transmission and emission spectra of giant exoplanets often display strong, well-resolved absorption/emission features spanning multiple spectral bands[30, 31, 32]. In such cases, the identity of major atmospheric constituents can often be inferred directly from visual inspection, with statistical model comparisons serving as confirmation rather than discovery[30]. These high-SNR observations highlight the unique strengths of JWST and have enabled confident chemical detections and precise abundance estimates in giant exoplanet atmospheres.
@@
More broadly, Bayesian model comparisons do not necessarily confirm the existence of a planetary archetype or physical process. An example is the atmospheric characterization of the sub-Earth-sized planet L 98-59 b[33], observed with JWST/NIRSpec from 2.7 to 5.2 m[34], Fig. 2. Prior to these observations, theoretical work[35] had identified L 98-59 b as a candidate for detecting extrasolar volcanism, predicting that SO2 absorption features could emerge if tidal heating rates resembled those of Io[36, 37]. The JWST data were found to be statistically consistent with a featureless spectrum based on analysis[34].
Nonetheless, Bayesian model comparisons favored atmospheric models over a bare-rock model at significance, and models including SO2 over models without it at , when performing nested model comparisons[34]. The study demonstrated a relative preference for models containing SO2 despite the lack of meaningful abundance constraints. Critically, it also acknowledged the limitations of this framework, emphasizing that model comparisons are conditional and that alternative scenarios, such as a no-atmosphere model, could not be ruled out. Confirming the existence of volcanic activity in an exoplanet would require complementary lines of evidence beyond the presence of SO2 in its atmosphere. Examples include reliable measurements of other chemical products resulting from volcanic activity as seen in the solar system[38] and even the detection of a circumstellar plasma tori[39], as previously acknowledged[35, 34].
@@ A Case Study: The MIRI/LRS Transmission Spectrum of K2-18b
We return to the case of K2-18 b to illustrate the challenges and risks of interpreting Bayesian model comparisons as detections of atmospheric gases in exoplanets. Here, we focus specifically on the JWST/MIRI transmission spectrum of K2-18 b[23] to examine methodology at the limits of what is currently observable with JWST; a prelude to the coming decades of exoplanet spectroscopy. We assess whether the observations are consistent with a featureless spectrum (see Methods). We also perform detailed atmospheric modeling as is routinely performed in the field, even for low signal-to-noise observations with weak spectral features[16, 34, 18, 20]. We emphasize that consistency with the null hypothesis of a flat spectrum does not preclude atmospheric interpretation, but rather provides critical context for assessing the robustness of any inferences drawn from the data. Likewise, the preference of an atmospheric model over a featureless/constant transit depth model in Bayesian comparison alone does not reject the null hypothesis. These considerations offer a key step towards determining whether spectral features are unambiguously present in data for reliable characterization of exoplanet atmospheres.
@@
We analyze the transmission spectrum with a range of atmospheric models with varying assumptions regarding the planetary thermo-physical-chemical conditions. These models range in complexity from fully parametric approaches that independently fit the chemical abundances and pressure–temperature structure without enforcing physical self-consistency (commonly referred to as ‘free retrievals’), to those that couple radiative–convective and thermochemical equilibrium with disequilibrium processes such as photochemistry and vertical mixing (see Methods). For example, equilibrium models may fail to capture atmospheres strongly influenced by disequilibrium chemistry, while fully parametric models may produce solutions that are not physically plausible[31].
@@
We conduct this analysis on the JWST/MIRI transmission spectrum of K2-18 b and find that all of our selected modeling frameworks provide an adequate fit to the data regardless of their underlying assumptions. We first consider physically motivated models that assume (i) radiative–convective–thermochemical equilibrium (RCTE), (ii) radiative–convective–photochemical equilibrium (RCPE), and (iii) chemical equilibrium. The model residuals, shown in Figure 3, are consistent with a normal distribution in all three cases. This indicates that the data do not reject the physics encapsulated within these assumptions (see Methods). Therefore, we conclude that the observations lack the constraining power necessary to distinguish between these physically distinct scenarios. Moreover, none of these models produce meaningful constraints on the atmospheric properties because the retrieved posterior distributions always closely reflect the input priors (see Methods). This result emphasizes that an adequate model fit does not necessarily imply that the inferred properties from a given model reflect the true atmospheric state of the planet. This is exacerbated when multiple models remain compatible with the data.
@@
We also consider a more flexible approach that allows the volume mixing ratios (i.e., abundances) of each chemical species to vary freely and independently. These models can attribute specific spectral features to particular chemical components. The Bayesian model comparison framework was originally introduced for this methodology[2]. Practitioners must decide which chemical species to include based on prior expectations of their plausibility in the atmosphere in question. This choice is also limited by both the number of species for which opacity data are available (several hundred[15]) and computational cost[40]. Here we include 21 chemical species motivated by theoretical calculations[41, 42], observations of Solar System planet atmospheres[43], and previous studies of K2-18 b[22] (see Methods). We include the proposed biosignature gases dimethyl sulphide (DMS) and dimethyl disulphide (DMDS), whose tentative presence has been reported in the atmosphere of K2-18 b[23]. We emphasize that this model implicitly assumes that no other gases beyond these 21 are present in the atmosphere of K2-18 b, and any inferences from this model must be interpreted within that context.
However, even with 21 species this model implicitly assumes that this small fraction of the complete prior hypothesis space (see e.g., Figure 1) is sufficient to explain the observed spectrum. As seen in Figure 4, several chemical species not considered in this 21 species model have cross-sections that overlap significantly in the wavelength range of these JWST/MIRI observations. Some of the prominent DMS and DMDS features, especially those at m and m, are similar for hydrocarbons not included in this 21 species model. Given these spectroscopic degeneracies and the low signal-to-noise ratio of these JWST/MIRI observations, multiple chemical species are likely to provide a reasonable fit and exacerbate the attribution problem. Even if many more species were included as free parameters, we would have to consider whether any meaningful, non degenerate constraints could be derived with such a large model and such limited data.
@@
These flexible retrievals do not yield conclusive results, consistent with the physically motivated models. Multiple gas combinations provide equally adequate fits to the observed spectrum. To quantify the support for individual chemical species, we compare the Bayesian evidence of the full model (including 21 chemical species) against that of a set of nested models, each neglecting one species at a time[2]. Among the 21 species considered, only five led to a change in Bayesian evidence corresponding to a model preference 1.2, with all significances well below 2 (Extended Data Table 1). The best-fit solutions from both the full model and the nested models yield in all cases which are indicative of overfitting. In the full model, the only species whose posterior deviated noticeably from the prior was propyne (C3H4), a hydrocarbon identified in the atmospheres of Titan[44], Saturn[45], Jupiter[46], Uranus[47], and Neptune[48].
@@
We now consider the reduction to a smaller model, as performed in the previous analysis of K2-18 b[23], to assess the corresponding model preferences. In their work, an atmospheric model with 32 parameters (20 molecular abundances) was initially considered and then reduced to a smaller model with 16 parameters (4 molecular abundances) including only methane (CH4), carbon dioxide (CO2), DMS and DMDS. The presence of the former two was informed by prior observations[22]. By comparing this four-gas model to a two-gas model excluding DMS and DMDS, the authors found a model preference between 2.9 and in favor of the inclusion of the sulfur species. We independently reconstruct this model configuration (see Methods) and recover a model preference of for the inclusion of DMS and DMDS relative to the model containing only CO2 and CH4.
We extend this analysis by constructing a series of models that include CH4, CO2 and a single additional chemical species. This expanded model comparison aims to test the robustness of the 3 confidence in the presence of DMS and DMDS. Specifically, we determine whether additional species that were not considered in the original analysis could equally or better explain the observations.
@@
We built 90 total models, each of which includes CH4, CO2, and a single hydrocarbon, covering all available cross-sections in the HITRAN database[14, 15]. We consider the presence of hydrocarbons for the following reasons: (i) CH4 is the dominant carbon carrier at low temperatures ( K)[41], and several hydrocarbons are natural product of CH4 photochemistry and thermochemistry at these temperatures (see Methods), (ii) several hydrocarbons have been detected or inferred to be present in the H-rich atmospheres of solar system bodies including Titan and Neptune[49, 48], and (iii) many hydrocarbons have prominent spectral features in the MIRI LRS wavelength range[50] (Figure 4).
@@
From this search, we find that every hydrocarbon increases the Bayesian evidence relative to the CH4+CO2-only model, with 59 models showing a preference greater than 2 (Extended Data Table 4). The best-performing model, shown in Figure 3, includes propyne (C3H4), yielding a 3.1 preference over the CH4+CO2-only model and a higher Bayesian evidence than the model containing DMS and DMDS. This result indicates that, even within the limited framework of comparing a given species or pair of species against a model with CH4+CO2 only, DMS and DMDS are not the most preferred gases. If we expand the model space further to include additional absorbers, we find that combining a small number of hydrocarbons with H2O can lead to scenarios where the inclusion of DMS and DMDS cause the Bayesian evidence to decrease, meaning that model comparison favors the absence of these species (see Methods).
@@
It is critical to note that the 3.1 preference for propyne is relative to a specific reference model and does not constitute a detection of propyne or any other species. This is analogous to the previous preference for DMS and DMDS. The apparently high detection significance of an individual gas can be misleading in small model spaces because there are fewer species available to explain the observations[2, 13]. When more gases are included, the degeneracy caused by overlapping absorption cross-sections is revealed (Figure 4), precluding unique attribution of features. As such, model preferences derived from limited model comparisons must be interpreted with caution, particularly when the overall constraining power of the data is weak, meaning that many molecules could likely explain any spectral features.

### v3

# Challenges in the detection of gases in exoplanet atmospheres

@@ Abstract
Claims of detections of gases in exoplanet atmospheres often rely on comparisons between models including and excluding specific chemical species. However, the space of molecular combinations available for model construction is vast and highly degenerate. Only a limited subset of these combinations is typically explored for any given detection. As a result, apparent detections of trace gases risk being artifacts of incomplete modeling rather than robust identification of atmospheric constituents, especially in the low signal-to-noise regime. Using the sub-Neptune K2-18 b as a case study, we show that recent biosignature claims vanish when the model space is expanded, with numerous alternatives providing equally good or better fits. We demonstrate that the significance of a claimed detection relies on the choice of models being compared, and that model preference does not in itself imply the presence of a specific gas. We recommend treating model comparisons instead as relative adequacy tests, which should be supported by theoretical predictions and complementary metrics of statistical significance in order to attribute a signal to a particular gas.
@@
While the atmospheric retrieval approach is generally regarded as agnostic and data-driven, it still has the potential to mislead regarding the strength of an apparent detection, and the subsequent interpretation of planetary conditions. The field has adopted Bayesian model comparison as a standard method to determine the presence of chemical species and physical processes in exoplanet atmospheres[2, 3, 4]. However, concerns have emerged that relying on the Bayesian evidence as the sole goodness-of-fit metric is insufficient for robust atmospheric interpretation[5]. Recent work has advocated for methods beyond the Bayesian evidence that provide insight into which parts of a dataset drive inferences[6, 7], and for averaging inferences over models instead of selecting between them[8]. However, the pitfalls of Bayesian model comparison in regimes lacking strong spectral features that can be clearly identified, attributed, and interpreted (see Key Criteria from ref.[9]) remain largely unexplored.
Here we demonstrate that Bayesian model comparison does not always translate into robust detections of atmospheric constituents. Pairwise model comparisons only quantify the relative fit to the data of one model against another. When all considered candidate models are poor representations of reality, the best-performing model is simply the least inadequate and may not necessarily lead to reliable interpretations of the data. Conversely, when all candidate models adequately fit a spectrum, a preference for one model over another does not rule out the worse-performing model. We demonstrate how these issues manifest in a case study of the sub-Neptune K2-18 b[10, 11], where the inferred atmospheric properties can be shaped as much by preconceived notions for what this planet ought to be like as by the observational data.
@@
Determining appropriate, data-driven model assumptions was an early challenge in analyzing exoplanet spectra given our limited a priori knowledge of these atmospheres. Bayesian model comparison was introduced to compare competing hypotheses in exoplanetary atmospheric studies[2]. Ref.[2] introduced an explicit methodology of computing the Bayesian evidence for one retrieval model that should cover the full prior hypothesis space, and comparing it to models for which one chemical species (or type of aerosol) was removed at a time from that otherwise full prior hypothesis space. Although exploring a complete prior hypothesis space may not be possible, as explained below, this remove-one-at-a-time approach was intended to ensure that a high confidence in the presence of a particular atmospheric constituent was reported only if no other reasonable constituent could have plausibly resulted in the observed data (Section 2.2.2 in ref.[2]). This approach was quickly adopted by the field, and became popular for its promise of applying “Occam’s razor” to penalize the fit of more complex models over simpler ones that were not sufficiently supported by the data.
@@
Given the practical impossibility of constructing a full prior hypothesis space, and no well-established conventions in the field, there is a risk of Bayesian model comparison being used to evaluate the relative preference between models even when neither adequately represents the full prior hypothesis space. Detections could then be reported based on Bayes factors between models that each include only a limited subset of chemical species, without comparison or contextualization relative to a more comprehensive hypothesis space. If two similarly incomplete or inadequate models are compared, there is a risk of incorrectly interpreting a model preference as a detection. The problem worsens if this model preference is converted into a “detection significance,” expressed in terms of [12], losing the relative nature of the metric and further contributing to misinterpretation.
@@
Use of model comparisons in isolation, without acknowledging their limitations, contributes to an opaque definition of “detection”. Any detection significance is computed relative to a chosen reference model and therefore strongly depends on the included species and parameters. Statements of model preference without qualification of the reference model are arbitrary. Furthermore, because the number of plausible chemical species is large — hundreds are available via the HITRAN database[13, 14] for example — meaningfully exploring any significant fraction of the hypothesis model space is computationally unfeasible. This practical limitation is an invitation to recognize the subjective nature in building an appropriate model space. Figure 1 shows how the number of unique models and pairwise model comparisons needed to exhaustively explore the detection space grows with the number of molecules considered. Even for a modest set of species, the number of unique compositional models for an atmosphere is , and the number of possible pairwise model comparisons is . While a complete exploration of model space is not currently practical, any detection claimed through pairwise model comparison should at minimum be contextualized within the broader set of plausible comparisons that could have yielded equal or greater significance. Credible detection claims demand systematic elimination of credible alternatives.
@@
The concept of detection in astronomy typically refers to the measurement of a signal above background noise, arising from detector electronics, intrinsic fluctuations in the photon flux, or other sources. Assuming normally distributed noise, the signal amplitude can be expressed in units of the noise’s standard deviation. A 3- detection, for example, implies that the observed signal exceeds the noise by three times its standard deviation. This is a fundamentally different measure than the Bayesian model comparisons explained above. However, the shared use of -confidence has resulted in ambiguities in the meaning of detection, and in some cases telescope resources are optimized to yield detections without clearly distinguishing whether the value is derived from a signal-to-noise argument or a model comparison.
In high signal-to-noise regimes, atmospheric detections rely on strong, well-resolved absorption/emission features spanning multiple spectral bands[15, 16, 17]. A clear example is the identification of CO2 in the atmosphere of WASP-39 b[15], where an absorption feature at 4.3m appears at significance, referring to the signal-to-noise ratio of the data; far stronger than previously predicted[18]. Attribution to CO2 is supported by atmospheric models for warm gas giants and expectations from thermochemical equilibrium[19, 20], and reinforced by the presence of additional CO2 features at other wavelengths, and across independent data reductions[21, 22]. Bayesian model comparisons confirm that models including CO2 best explain the observed features, but in this context, the comparison does not constitute the detection itself; it merely affirms the attribution.
While CO2 in WASP-39 b demonstrates a clear detection grounded in strong data and physical plausibility, many recent JWST spectra of smaller exoplanets operate near the limits of low signal-to-noise, where interpretation becomes more ambiguous. JWST has delivered spectra of some of the smallest and coldest exoplanets studied to date[23, 24, 25, 26, 27, 28, 29, 30, 31], yet several of these observations remain statistically consistent with a featureless spectrum (see Figure 2). Testing whether a spectrum is consistent with an absence of features is often conducted using a frequentist approach, which tests whether a null hypothesis can be rejected at a given confidence level (see Methods and Extended Data Figure 1). This contrasts with Bayesian model comparison, which quantifies the relative preference between two models.
@@
An example of this distinction is the atmospheric characterization of the sub-Earth-sized planet L 98-59 b[32], observed with JWST/NIRSpec[33], Figure 2. Prior to these observations, theoretical work[34] had identified L 98-59 b as a candidate for detecting extrasolar volcanism, predicting that SO2 absorption features could emerge if tidal heating rates resembled those of Io[35, 36]. The JWST data were found to be statistically consistent with a featureless spectrum based on analysis[33]. Nonetheless, Bayesian model comparisons favored atmospheric models over a bare-rock model at significance, and models including SO2 over models without it at [33]. Critically, the study acknowledged the limitations of this framework, emphasizing that model comparisons are conditional and that a no-atmosphere model could not be ruled out. Confirming the existence of volcanic activity in an exoplanet would require complementary lines of evidence beyond the presence of SO2 in its atmosphere, such as measurements of other chemical products resulting from volcanic activity[37] or the detection of a circumstellar plasma torus[38].
Degeneracies between atmospheric parameters further contribute to the complexity of exoplanet atmospheric retrieval. Numerous molecules have overlapping absorption bands, meaning that several chemical species can often explain a given feature in an observed spectrum. An example of this degeneracy appears in the analysis of the –m HST/WFC3 transmission spectrum of the sub-Neptune K2-18 b[10]. The spectrum exhibits a prominent feature near m, initially attributed to H2O at [10], [39], and [40, 41] significance by four independent analyses, all using Bayesian model comparisons. Interpreted as a frequentist significance, this would imply a 99.7% probability that H2O is present. However, a subsequent study demonstrated that model realizations with methane as the dominant absorber instead of water also provide adequate fits to the data ()[42].
Follow-up observations with JWST/NIRISS and NIRSpec, spanning a broader spectral range (0.9–5.2m), revealed a preference for methane at – significance, depending on model assumptions, with no strong evidence for H2O[29, 43]. The extended coverage from JWST helped resolve the degeneracy between H2O and CH4 at HST/WFC3 wavelengths. Comparing the HST and JWST observations, it appears that some previous attributions to H2O may have been influenced by two data points near the blue edge of the WFC3 bandpass (see Figure 6 in ref.[29]). Furthermore, one earlier study did not consider CH4 absorption in its atmospheric models[39]. The case of the HST/WFC3 spectrum of K2-18 b highlights the risks of interpreting model preference as a definitive chemical detection—particularly when the data are limited in spectral coverage and models are highly simplified.
@@ A Case Study: The MIRI/LRS Transmission Spectrum of K2-18 b
We further explore the case of K2-18 b, now focusing on the JWST/MIRI transmission spectrum[30], to illustrate the risks of interpreting Bayesian model comparisons as gas detections. We find that the observations are consistent with a featureless spectrum, with any apparent signals being indistinguishable from statistical noise (see Methods, Extended Data Figure 1). We also perform detailed atmospheric modeling as is routinely performed in the field, even for low signal-to-noise observations with weak spectral features[23, 33, 25, 27]. We emphasize that consistency with the null hypothesis of a flat spectrum (see Extended Data Figure 2) does not preclude atmospheric interpretation, but rather provides critical context for assessing the robustness of any inferences drawn from the data. Likewise, the preference of an atmospheric model over a featureless/constant transit depth model in Bayesian comparison alone does not reject the null hypothesis.
@@
We analyze the transmission spectrum with atmospheric models ranging in complexity from fully parametric approaches that independently fit the chemical abundances and pressure–temperature structure without enforcing physical self-consistency (commonly referred to as ‘free retrievals’), to those that couple radiative–convective and thermochemical equilibrium with disequilibrium processes such as photochemistry and vertical mixing (see Methods, Extended Data Figure 3). Each approach has limitations: for example, equilibrium models may fail to capture atmospheres strongly influenced by disequilibrium chemistry, while fully parametric models may produce solutions that are not physically plausible[16].
@@
We find that each modeling framework provides an adequate fit to the spectrum regardless of its underlying assumptions (see Methods, Extended Data Figure 4). We first consider physically motivated models that assume (i) radiative–convective–thermochemical equilibrium (RCTE), (ii) radiative–convective–photochemical equilibrium (RCPE), and (iii) chemical equilibrium. The model residuals, shown in Figure 3, are consistent with a normal distribution in all three cases. This indicates that the data do not reject the physics encapsulated within these assumptions. We conclude that the observations lack the constraining power necessary to distinguish between these physically distinct scenarios. Moreover, no model obtains meaningful constraints on the atmospheric properties, with posterior distributions closely reflecting the input priors. This emphasizes that an adequate model fit does not necessarily imply that the inferred properties from a given model reflect the true atmospheric state of the planet, particularly when multiple models remain compatible with the data.
@@
We also consider a more flexible approach that allows the volume mixing ratios (i.e., abundances) of each chemical species to vary freely and independently. These models can attribute specific spectral features to particular chemical components. The Bayesian model comparison framework was originally introduced for this methodology[2]. Practitioners must decide which chemical species to include based on prior expectations of their plausibility in the atmosphere in question. This choice is limited by both the number of species for which opacity data are available (several hundred[14]) and computational cost[44]. Here we include 21 chemical species motivated by theoretical calculations[20, 45], observations of Solar System planet atmospheres[46], and previous studies of K2-18 b[29] (see Methods, Extended Data Figure 3). We include the proposed biosignature gases dimethyl sulfide (DMS) and dimethyl disulfide (DMDS), whose tentative presence has been reported in the atmosphere of K2-18 b[30].
@@
These flexible retrievals of the K2-18 b MIRI observations do not yield conclusive results, consistent with our findings with the physically motivated models. Multiple gas combinations provide adequate fits to the observed spectrum. To quantify the support for individual chemical species, we compare the Bayesian evidence of the full model (including 21 chemical species) against that of a set of nested models, each neglecting one species at a time[2]. Among the 21 species considered, only five led to a change in Bayesian evidence corresponding to a model preference 1.2 () for JExoRES and only two for JexoPipe, with all significances well below 2, including CH4, CO2, DMS, and DMDS (Extended Data Table 1). The best-fit solutions from both the full model and the nested models yield for JExoRES and for JexoPipe, which are indicative of overfitting in all cases. In the full model, the only species whose posterior deviated noticeably from the prior for both the JExoRES and Jexopipe reductions was propyne (C3H4), a hydrocarbon identified in the atmospheres of Titan[47], Saturn[48], Jupiter[49], Uranus[50], and Neptune[51].
We emphasize that this model implicitly assumes that no other gases beyond these 21 are present in the atmosphere of K2-18 b. Any inferences from this model must be interpreted within that context. As seen in Figure 4, several chemical species not considered in this 21-species model have cross-sections that overlap significantly in the wavelength range of these JWST/MIRI observations. Some of the prominent DMS and DMDS features, especially those at m and m, are similar for hydrocarbons not included in this 21 species model. Given these spectroscopic degeneracies and the low signal-to-noise ratio of these JWST/MIRI observations, multiple chemical species are likely to provide a reasonable fit and exacerbate the attribution problem. Indeed, we find that expanding this model to include 29 absorbers yields similar results (see Methods).
@@
We now consider the reduction to a smaller model, as performed in the previous analysis of K2-18 b[30], to assess corresponding model preferences. In that work, an atmospheric model with 32 parameters (20 molecular abundances) was initially considered and then reduced to a smaller model with 16 parameters (4 molecular abundances) including only CH4, CO2, DMS and DMDS. The presence of CH4 and CO2 was informed by prior observations[29]. By comparing this four-gas model to a two-gas model excluding DMS and DMDS, the authors found a model preference between 2.9 and in favor of the inclusion of the sulfur species. We independently reconstruct this model configuration (see Methods, Extended Data Table 2) and recover a model preference of for the inclusion of DMS or DMDS relative to the model containing only CO2 and CH4 — a model comparison approach that deviates from standard practice of comparing nested models (see Methods, Extended Data Table 3, Extended Data Figure 5).
@@
We extend this analysis with the aim of testing the robustness of the 3 confidence in the presence of DMS and DMDS. Specifically, we determine whether additional species that were not considered in the original analysis could equally or better explain the observations. We built 90 total models, each of which includes CH4, CO2, and a single hydrocarbon, covering all available HITRAN cross-sections[13, 14]. We consider hydrocarbons for the following reasons: (i) CH4 is the dominant carbon carrier at low temperatures ( K)[20], and several hydrocarbons are products of CH4 photochemistry and thermochemistry at these temperatures (see Methods, Extended Data Figure 3), (ii) several hydrocarbons have been identified in the H-rich atmospheres of solar system bodies including Titan and Neptune[52, 51], and (iii) many hydrocarbons have spectral features in the MIRI/LRS wavelength range[53] (Figure 4).
@@
We find that almost every hydrocarbon increases the Bayesian evidence relative to the CH4+CO2-only model, with 60 models showing a preference greater than 2 for at least one reduction of the data (Extended Data Table 4 and Extended Data Table 5). The best-performing model, shown in Figure 3, includes propyne (C3H4), yielding a preference over the CH4+CO2-only model and a higher Bayesian evidence than the model including DMS and DMDS for both reductions. This indicates that, even within the limited framework of comparing a given species or pair of species against a model with CH4+CO2 only, DMS and DMDS are not the most preferred gases. By expanding the model space to include additional absorbers, we find that combining a small number of hydrocarbons with H2O can lead to scenarios where the inclusion of DMS and DMDS cause the Bayesian evidence to decrease, meaning that model comparison favors the absence of these species (see Methods).
@@
It is critical to note that the 3.1 preference for propyne is relative to a specific reference model, and does not constitute a detection of propyne (or any other species). This is analogous to the previous preference for DMS and DMDS. The apparently high detection significance of an individual gas is misleading in small model spaces where fewer species are available to explain the observations[2, 12]. When more gases are included, degeneracies between overlapping absorption cross-sections preclude unique attribution of features (Figure 4). Model preferences derived from limited comparisons must be interpreted with extreme caution, particularly when the overall constraining power of the data is weak.